OpenStudy (anonymous):

What is tan A for the triangle shown? A. 4/5 B. 3/4 C. 5/4 D. 3/5

OpenStudy (anonymous):

Hers is the Triangle.

OpenStudy (anonymous):

3/4 because tan means adjacent over opposite. so 12/16= 6/8= 3/4

OpenStudy (anonymous):

Thank Yuo so much! I'm posting some mmore can you help with them too cause' i've been stuck on them all day.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Find the value of x to the nearest tenth. A. 3.7 B. 11.5 C. 3.3 D. 4.5

OpenStudy (anonymous):

2. What is the measure of angle A in the triangle below? A. 71.1 B. 18.9 C. 29.2 D. 90

OpenStudy (anonymous):

for the second problem, if tan is opposite to adjacent it means that tan42=x/5. so isolate x and it will give you the answer.

OpenStudy (anonymous):

What is the side length b in the triangle below? A. 23 B. 16.6 C. 46 D. 32.8

OpenStudy (anonymous):

again, for the third problem you need to imput the values. so tan A=opp/adj=12/35. Now, to find A you will need your calculator. A= arctant 12/35. On your calculator, arctan is tan^(-1)

OpenStudy (anonymous):

you need to remember your identities. cos=adj/hyp sin=opp/hyp tan=opp/adj your last problem deels with one of these three identities. Try to find the answer. The idea is the same as the other problems.

OpenStudy (anonymous):

Oh ok, thank you so much that makes more sense:)